How many radial nodes and how many angular nodes does each of the following orbitals have?

4d

6s

4f
Answer #1

Number of angular nodes = l

Number of radial nodes = n-l-1

Total node = n-1

a)

For 4d orbital, n=4 and l=2

So,

Number of radial nodes = n-l-1

= 4 - 2 - 1

= 1

Number of angular nodes = l

= 2

Answer: 1, 2

b)

For 6s orbital, n=6 and l=0

So,

Number of radial nodes = n-l-1

= 6 - 0 - 1

= 5

Number of angular nodes = l

= 0

Answer: 5, 0

c)

For 4f orbital, n=4 and l=3

So,

Number of radial nodes = n-l-1

= 4 - 3 - 1

= 0

Number of angular nodes = l

= 3

Answer: 0, 3